OpenStudy (bbb911):

This is Geometry: solve the system using substitution or elimination: 2x-4y=-2 4x+8y=0

OpenStudy (anonymous):

Heck ya! I love these so much. lul

OpenStudy (anonymous):

Equation one: 2x-4y=-2 Equation two: 4x+8y=0 Notice that the signs on the two Ys are different. So they will cancel out well once we make the coefficients the same. To do this, multiple the top equation by two. Equation one turns into: 2(2x-4y=-2) Or... 4x-8y = -4 Note that all I did was double each value. So let's call this 'equation three' and temporarily replace equation one with it. Equation 3 = 4x-8y = -4 Equation 2 = 4x+8y = 0 Add them up and you get equation 4: 8x + 0 = -4 Divide by 8: x = -4/8 -4/8 turns into -1/2 after simplifying. So... we know X is negative one half.

OpenStudy (anonymous):

Take equation one: 2x-4y=-2 Plug in x as -1/2 2(-1/2) - 4y = -2 -1 - 4y = -2 -4y = -1 -y = -1/4 multiply by negative 1: y= 1/4

OpenStudy (anonymous):

Hmm... same answer as before... So, we have x = -1/2 and y = 1/4 Let's test these one more time. 2x-4y=-2 2(-1/2)-4(1/4) = -2 -1 - 1 = -2 -2 = -2 (TRUE!) 4x+8y=0 4(-1/2) + 8(1/4) = 0 -2 + 2 = 0 0 = 0 (TRUE) Guess I was in fact right the first time. Lol.

OpenStudy (bbb911):

thhank you so much

OpenStudy (anonymous):

No problemo! BTW, do let me know if any of these don't make sense. I can definitely assist with explaining any weird steps.

OpenStudy (bbb911):

oh ok sure thing